Global Services is considering a promotional campaign that will increase annual credit sales by $610,000. The company will require investments in accounts receivable, inventory, and plant and equipment. The turnover for each is as follows: Accounts receivable 4 times 4 times Plant and equipment 2 times All $610,000 of the sales will be collectible. However, collection costs will be 2 percent of sales, and production and selling costs will be 70 percent of sales. The cost to carry inventory will be 10 percent of inventory. Depreciation expense on plant and equipment will be 10 percent of plant and equipment. The tax rate is 30 percent. a. Compute the investments in accounts receivable, inventory, and plant and equipment based on the turnover ratios. Add the three together Accounts receivable Inven Plant and UI Total Investment$ b. Compute the accounts receivable collection costs and production and selling costs and then add the two figures together Collection cost Production and selling costs Total collection, production, ands selling costs 0 c. Compute the costs of carrying inventory Cost of carrying inv d. Compute the depreciation expense on new plant and equipment. Depreciation e. Compute the total of all costs from parts b through d. Total costs f. Compute income after taxes. after taxes g-1. What is the aftertax rate of return? (Input your answer as a percent rounded to 2 decimal places.) Aftertax rate o return
Answer #1
A B C D E F G H I
2
3 Increase in Annual Sales $610,000
4
5 Accounts receivable turnover 4 times
6 Inventory turnover 4 times
7 Plant and equipment 2 times
8
9 a)
10
11 Accounts receivable turnover can be calculated as follows:
12 Accounts receivable turnover = Credit Sales / Average Accounts Receivable
13
14 Investment in Accounts receivable can be calculated as follows:
15 Investment in Accounts receivable =Increase in Sales / Accounts receivable turnover
16 $152,500 =D3/D5
17
18 Similarly,
19 Investment in Inventory $152,500 =D3/D6
20 Investment in Plant and Equip. $305,000 =D3/D7
21
22 Hence,
23 Accounts Receivable $152,500
24 Inventory $152,500
25 Plant & Equipment $305,000
26 Total investment $610,000 =SUM(D23:D25)
27
28 b)
29
30 Collection cost 2% of sales
31 Production and selling costs 70% of Sales
32
33 Collection cost $12,200 =D3*D30
34 Production and selling costs $427,000 =D3*D31
35 Total Collection, production and selling costs $439,200 =SUM(D33:D34)
36
37 c)
38
39 Cost of carrying inventory 10% of inventory
40
41 Cost of carrying inventory $15,250 =D24*D39
42
43 d)
44
45 Depreciation expense 10% of plant and equipment
46
47 Depreciation expense $30,500 =D25*D45
48
49 e)
50
51 Total cost $484,950 =D35+D41+D47
52
53 f)
54
55 Increase in Sales $610,000
56 Total Expense $484,950
57 Income before tax $125,050 =D55-D56
58 Tax Expense (30%) $37,515 =D57*30%
59 After-tax Income $87,535 =D57-D58
60
61 Hence after tax income is $87,535
62
63 g)
64
65 After-tax rate of return =After tax income / Total investment
66 14.35% =D61/D26
67
68 After-tax rate of return 14.35%
69
70